Express the confidence interval ( 84.1 % , 91.3 % ) ( 84.1 % , 91.3 % ) in the form of ˆ p ± M E p ^ ± M E .
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.841+0.913) / 2 =0.877

Margin of error = ME = Upper confidence interval - = 0.913 - 0.877=0.036

In the form of :

- ME

0.877 +/-0.036

87.7%+/-3.6%